OpenStudy (annie96):

3. Andrea wants to deposit money into a bank account that earns 2.5% simple interest. Use the formula to find the amount of money that she should deposit so that she earns $2500 after 4.5 years. Show and explain your work.

OpenStudy (madhu.mukherjee.946):

A = P(1 + rt)

OpenStudy (madhu.mukherjee.946):

First, converting R percent to r a decimal r = R/100 = 2.5%/100 = 0.025 per year, then, solving our equation A = 2500(1 + (0.025 × 4.5)) = 2781.25 A = $ 2,781.25 The total amount accrued, principal plus interest, from simple interest on a principal of $ 2,500.00 at a rate of 2.5% per year for 4.5 years is $ 2,781.25.

OpenStudy (annie96):

thank you.